(1 point) Find the length of the curve defined by

y=18(8x2−1ln(x))y=18(8x2−1ln⁡(x))

from x=4x=4 to x=8

(1 point) Find the area of the region enclosed by the curves:
2y=4x−−√,y=4,2y=4x,y=4, and 2y+1x=52y+1x=5

HINT: Sketch the region!

y=2+1/x4,y=2,x=4,x=9;y=2+1/x4,y=2,x=4,x=9;
about the x-axis.

(1 point) Find the length of the curve defined by y = $(8x? – 1 In(x)) from x = 4 to x = 8(1 point) Find the volume of the solid obtained by rotating the region bounded by the given curves about the specified axis.(1 point) Find the area of the region enclosed by the curves: 2y = 4VX, y = 4, and 2y + 1x = 5 HINT: Sketch the region!
Answer #1

given by Length of a Curve y=fln) from xaa do xab is 6 dy Aast to x=B inquestion. Nah texas se fa=4&h=8 & Hex) = f (8x7 - las५- 2 + 15-20५ १०१ about -axis y = frái) = 52.81) - 2 14 (64), orde[4] then volume of the solid formed by revolving the region